• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Punjabi

    Log entry time 00:09:29 on June10,2000

    Entry number 43297

    keyword=Shift Summary

    Started shift with run no. 1847

    16:00 to 17:25 run 1847 to 1851

    lost beam at 17:25 , Cavity problem

    17:40 asked for access to fix problem of hole in x-trajectory
    spectra in hadron arm. Suspicion is that one of the scintillater paddle in
    s1 plane is tilted.

    18:30 scintillator s1-6 was rotated to move VDC in. They forgot to put it back in the correct position. All runs from 1823 to 1851 were with this paddle in the wrong position.

    18:55 beam back , started run 1852
    19:05 beam off for 5 minuites
    19:10 beam back, continue with run 1852
    19:25 stop run 1852
    19:26 start run 1853
    19: 49 stop run 1853

    Opened collimator in both spectrometer

    20:06 started run 1855 . First run with open collimator in both spectrometer.
    20:15 stopped run 1855

    changed prescale factors because event rate increased due to open collinators.

    20:16 started run 1856 with new prescale factors
    20:26 stopped run 1856

    no beam for about 10 minuites

    20:27 started run 1857
    20:47 stopped run 1857
    20:47 start run 1858
    21:00 stopped run 1858
    21:01 started run 1859
    21:14 stopped run 1859
    21:15 started run 1860
    21:27 stopped run 1860
    21:28 started run 1861
    21:41 stopped run 1861
    21:42 started run 1862
    21:47 stopped run 1862

    run 1862 is not a good run

    Noticed that event rate jumped sudenly. In trying to find out the reason for this found that helium level in electron Q3 had dropped to 44 % and so there was a quench in this quad. Asked for an access to reset the quench .

    22:25 beam back after resetting quad 3 in elrctron arm.
    22:25 start run 1863
    22:41 stopped run 1863
    22:42 start run 1864
    22:54 stopped run 1864
    22:55 start run 1865
    23:07 stop run 1865
    23:07 start run 1866
    23:19 stop run 1866
    23:19 start run 1867
    23:34 stop run 1867

    23:35 Helium level dropped again in electron quad 3 . So asked for access to reset the quench.

    23:55 Zhengwei went to hall to reset the quench.

    23:55 Called Mark stevens (Tech on call). He said may be some connections are bad and that is what is causing the problem. He thinks that Q3 quenches first and then helium level drops. We do not know at this time which one happens first, quenching or drop in helium level. He said if problem occurs again, page him and he will come in.

    00:10 Quad 3 is reset and it is comming up. We will keep taking data for now until next time the same problem occur.